Find two numbers whose difference is 10 and1/6 of whose sum is 11.
Answer by nabla(475)

a-b=10
(1/6)(a+b)=11
Solve by substitution.
From the first equation: a=10+b
Put that into the second equation.
(1/6)(10+2b)=11
10+2b=66
2b=56
b=28
Now from the first equation,
a-28=10
a=38.
So, 28 and 38 are two such numbers.